Session: Nov 2010 (session 3)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (scalambra at Thu Nov 04 05:07:37 2010) Success:

Torun (Kaz Borkowski at Thu Nov 04 06:33:17 2010) Success:
No known problems.

Effelsberg (Uwe Bach at Fri Nov 05 11:40:41 2010) Minor failures: ANT
Oil leakage at the elevation gear. Had to stop from 7:40 to 14 UT for repair.

Effelsberg (Uwe Bach at Fri Nov 05 11:41:59 2010) Success:
Sorry, worng exp. This one was okay, but EL030C stopped because of oil leakage.

Westerbork (Geert Kuper at Fri Nov 05 19:32:31 2010) Success:
Apart from RFI no known problems.

Onsala (Michael Lindqvist at Sat Nov 06 08:31:15 2010) Failure:
Could not observe, to much wind.